A family has 6 children. Assume that each child is as likely to be
a boy as it is to be a girl. Find the probability that the family
has 6 girls if it is known the family has at least one girl.

Answer by Edwin McCravy(20054)

The complement event of at least one girl is the event of 6 boys.

The probability of no girls, or all boys, is:

Probability of a boy 1st, and a boy 2nd, and a boy 3rd,
and a boy 4th, and a boy 5th, and a boy 6th. 

(1/2)(1/2)(1/2)(1/2)(1/2)(1/2) = (1/2)6 = 1/64.

So the probability of at least 1 girl is the complement
event, whose probability is:

1 - 1/64 = 63/64.

Edwin